OpenStudy (anonymous):

At what speed does a 1900 kg compact car have the same kinetic energy as a 1.7 x 10^4 kg truck going 25 km/hr?

OpenStudy (anonymous):

it has to be in km/hr I'm stuck on this one please help!

OpenStudy (anonymous):

E = 1/2 * m * v^2 E_truck = 0.5 * 1.7 * 10^4 [kg] * 25 [km/h] calculate and set E_truck equal to E_car ... and calculate your speed (v) from that

OpenStudy (anonymous):

oops E_truck = 0.5 * 1.7 * 10^4 [kg] * 25 [km/h] * 25 [km/h] , of course ;>

OpenStudy (anonymous):

mass of truck=1.7 x 10^4 kg velocity of the truck=25km/hr=6.944 m/s E(truck) = 1/2 * m * v^2=0.5*1.7*10^4*(6.944^2) =4.099 × 10^5kgm^2/s^2 E(truck)=E(car) Mass of compact car=1900 kg E (car)=0.5*1900*V^2 0.5*1900*V^2=4.099 × 10^5 \[V^2= (\frac{4.099 × 10^5}{0.5*1900})\] V=20.77m/s=74.772 km/h Speed = approx 74.772 km/h